# Single-file dynamics with general charge measures

### v2

@@ C.1.1 Saddle point on the imaginary axis
    which is manifestly real and positive for $x\in\mathbb{R}$. Suppressing inessential indices and setting $k_{\pm}={\rm i}x$, $x\in\mathbb{R}$, the saddle point equations can be written as
@@
    $h^{\prime}(x)$ $=\frac{r^{\prime\prime}(x)r(x)-[r^{\prime}(x)]^{2}}{[r(x)]^{2}}=\frac{\iint_{\mathbb{R}^{2}}{\rm d}^{2}\omega_{\mu}\left(c_{1}^{2}-c_{1}c_{2}\right)e^{x(c_{1}+c_{2})}}{[r(x)]^{2}}$ (107)
    $=\frac{\iint_{\mathbb{R}^{2}}{\rm d}^{2}\omega_{\mu}\left(c_{1}^{2}+c_{2}^{2}-2c_{1}c_{2}\right)e^{x(c_{1}+c_{2})}}{2[r(x)]^{2}}$ (108)
@@
    we find that $h$ is continuous and strictly monotonous, with equality in (109) being achieved only for a trivial measure supported on a single point. We also observe that
    $h(x)=\frac{r^{\prime}(x)}{r(x)}=\frac{\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,ce^{cx}}{r(x)}=\frac{\int_{-\infty}^{0}{\rm d}\omega_{\mu}\,ce^{cx}+\int_{0}^{\infty}{\rm d}\omega_{\mu}\,ce^{cx}}{r(x)}.$ (110)
    Both terms in the numerator are non-zero and of different sign. Moreover, they can be made arbitrarily large or small by sending $x\to\pm\infty$. If $\omega_{\mu}$ is finitely supported, the integrals in (110) localize around the support’s boundary points
@@
    which can be understood as taking the limit $c_{\pm}\to\pm\infty$. Returning to Eq. (106), we note that the expression on the left hand side changes sign for $x\to\pm\infty$ as a a consequence of (111) and bounds on the charge current (46-49). The same conclusion holds in the case of unbounded support via (112).
    Since $h$ is continuous and strictly monotonic, it follows that Eq. (106) has exactly one solution $x_{0}\in\mathbb{R}$. Reinstating the notation of Eqs. (104), we denote these solutions on the imaginary axis as
@@
    C.1.2 Saddle point localization
    Expanding the function $f_{\pm}$ around the saddle points ${\rm i}\kappa_{\pm}$ we find
@@ C.1.2 Saddle point localization
    Importantly, the coefficients at quadratic order are strictly negative (for $j_{p}\neq 0$) by virtue of (105),(109) and the saddle points are non-degenerate. Now consider the contours of steepest descent $\mathcal{I}_{\pm}$ across the saddle points ${\rm i}\kappa_{\pm}$ that are locally parallel to the real axis
    ±=$\mathcal{I}_{\pm}=\{k\in\mathbb{C}\,|\,{\rm Im}\,[f_{\pm}(k)]={\rm const},\enspace\kappa_{\pm}\in\mathcal{I}_{\pm},\enspace\mathcal{I}_{\pm}\parallel\mathbb{R}\textrm{ at }\kappa_{\pm}\}.$ (115)
@@
    $I_{c|p}^{\pm}(j_{c}|j_{p})=-\lim_{t\to\infty}t^{-\alpha}\log\int_{\mathcal{I}_{\pm}}\frac{{\rm d}k}{2\pi}e^{t^{\alpha}f_{\pm}(k)}=-f_{\pm}({\rm i}\kappa_{\pm})=\pm\left[j_{p}\log\hat{\omega}_{\mu_{\pm}}({\rm i}k_{\pm})-k_{\pm}j_{c}\right].$ (116)
    C.2 Generic properties of the split rate functions
    We now prove the properties (1-4), see Eqs. (61-64) of the main text.
    1. (1)
      Claim: $I_{c|p}^{\pm}(j_{c}|j_{p})$ vanish as $j_{p}\to 0^{\mp}$ for $j_{c}=0$.
      Proof: For $j_{c}=0$, the saddle point equation (60) reduces to $\partial_{k}\hat{\omega}_{\mu_{\pm}}(k)|_{k={\rm i}\kappa_{\pm}}=0$, showing that $\kappa_{\pm}$ is independent of $j_{p}$. The result now follows directly from taking the limit $j_{p}\to 0^{\mp}$ in (59).
    2. (2)
      Claim: Derivatives of $I_{c|p}^{\pm}(j_{c}|j_{p})$ with respect to $j_{p}$ are finite and sign-definite for $j_{c}=0$.
@@ C.2 Generic properties of the split rate functions
      ±$\pm\partial_{j_{p}}I_{c|p}^{\pm}(j_{c}|j_{p})=\log\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm}),$ (117)
      where reality of $I_{c|p}^{\pm}$ follows from (100). It is straightforward to show that $\hat{\omega}_{\mu_{\pm}}$ is strictly convex on the imaginary axis, $\partial_{x}^{2}\hat{\omega}_{\mu_{\pm}}({\rm i}x)=\int_{-\infty}^{\infty}{\rm d}\omega_{\mu_{\pm}}\,c^{2}e^{cx}>0$ while for $j_{c}=0$, the saddle point equation specifies that the derivative of $\hat{\omega}_{\mu_{\pm}}$ vanishes. Since the saddle points ${\rm i}\kappa_{\pm}$ are on the imaginary axis they are precisely the minima of $\hat{\omega}_{\mu_{\pm}}$ on the imaginary axis. Taking the value at the origin $\hat{\omega}_{\mu_{\pm}}(0)=1$, we have
@@
      where finiteness follows from positivity of $\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})$. Plugging (118) into (117) gives the result (62).
    3. (3)
      Claim: Derivatives of $I_{c|p}^{\pm}(j_{c}|j_{p})$ with respect to $j_{p}$ diverge as $j_{p}\to j_{p}^{\mp}$ for $j_{c}\neq 0$.
@@
      It immediately follows that $\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})=\int_{-\infty}^{\infty}{\rm d}\omega_{\mu_{\pm}}(c)\,e^{\kappa_{\pm}c}\to\infty$ as $|\kappa_{\pm}|\to\infty$, which together with (117) gives the result (63).
    4. (4)
      Claim: The Hessian determinant of $I_{c,p}^{\pm}(j_{c},j_{p})$ is positive.
@@
      $\frac{{\rm d}}{{\rm d}u}\kappa_{\pm}(u)=-\frac{\omega_{\mu_{\pm}}^{2}({\rm i}\kappa_{\pm})}{\partial_{k}^{2}\hat{\omega}_{\mu_{\pm}}(k)|_{k={\rm i}\kappa_{\pm}}\omega_{\mu_{\pm}}({\rm i}\kappa_{\pm})+\left[\partial_{k}\hat{\omega}_{\mu_{\pm}}(k)|_{k={\rm i}\kappa_{\pm}}\right]^{2}}=\frac{1}{h^{\prime}(\kappa_{\pm})}>0,$ (124)
@@
      with positivity a consequence of (124) and strict convexity of $I_{p}$, giving the desired result (64).

### v3

@@ D.1.1 Saddle point on the imaginary axis
is well-defined and is manifestly real and positive for $x\in\mathbb{R}$. Suppressing inessential indices and setting $k_{\pm}={\rm i}x$, $x\in\mathbb{R}$, the saddle point equations can be written as
@@
$h^{\prime}(x)$ $=\frac{r^{\prime\prime}(x)r(x)-[r^{\prime}(x)]^{2}}{[r(x)]^{2}}$
$=\frac{\iint_{\mathbb{R}^{2}}{\rm d}^{2}\omega_{\mu}\left(c_{1}^{2}-c_{1}c_{2}\right)e^{x(c_{1}+c_{2})}}{[r(x)]^{2}}$
$=\frac{\iint_{\mathbb{R}^{2}}{\rm d}^{2}\omega_{\mu}\left(c_{1}^{2}+c_{2}^{2}-2c_{1}c_{2}\right)e^{x(c_{1}+c_{2})}}{2[r(x)]^{2}}$
@@
we find that $h$ is continuous and strictly monotonous, with equality in (93) being achieved only for a trivial measure supported on a single point. We also have
$h(x)$ $=\frac{r^{\prime}(x)}{r(x)}=\frac{\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,ce^{cx}}{r(x)}$
$=\frac{\int_{-\infty}^{0}{\rm d}\omega_{\mu}\,ce^{cx}+\int_{0}^{\infty}{\rm d}\omega_{\mu}\,ce^{cx}}{r(x)}.$ (94)
Both terms in the numerator of (94) are non-zero and of different sign. Moreover, they can be made arbitrarily large or small by sending $x\to\pm\infty$. If $\omega_{\mu}$ is finitely supported, the integrals in (94) localize around the support’s boundary points
@@
which can be understood as taking the limit $c_{\pm}\to\pm\infty$. Returning to Eq. (92), we note that the expression on the left hand side changes sign for $x\to\pm\infty$ as a a consequence of (95) and bounds on the charge current (80-83). The same conclusion holds in the case of unbounded support via (96). Since $h$ is continuous and strictly monotonic, it follows that Eq. (92) has exactly one solution $x_{0}\in\mathbb{R}$. Reinstating the notation of Eqs. (90), we denote these solutions on the imaginary axis as
@@
D.1.2 Saddle point localization
Around the saddle points ${\rm i}\kappa_{\pm}$ the functions $f_{\pm}$ read
@@ D.1.2 Saddle point localization
Importantly, the coefficients at quadratic order are strictly negative (for $j_{p}\neq 0$) by virtue of (91) and (93) so that the saddle points are non-degenerate. Now consider the contours of steepest descent $\mathcal{I}_{\pm}$ across the saddle points ${\rm i}\kappa_{\pm}$ that are locally parallel to the real axis
±=$\mathcal{I}_{\pm}=\{k\in\mathbb{C}\,|\,{\rm Im}\,[f_{\pm}(k)]={\rm const},\enspace\mathcal{I}_{\pm}|_{\kappa_{\pm}}\parallel\mathbb{R}\}.$ (99)
@@
$I_{c|p}^{\pm}(j_{c}|j_{p})$ $=-\lim_{t\to\infty}t^{-\alpha}\log\int_{\mathcal{I}_{\pm}}\frac{{\rm d}k}{2\pi}e^{t^{\alpha}f_{\pm}(k)}$ (100)
$=-f_{\pm}({\rm i}\kappa_{\pm})=\pm\left[j_{p}\log\hat{\omega}_{\mu_{\pm}}({\rm i}k_{\pm})-k_{\pm}j_{c}\right].$
D.2 Generic properties of the split rate functions
We now prove Equations 42, 43, 44 and 45, reproduced below for convenience.
Claim 1: $I_{c,p}^{\pm}(j_{c}=0,j_{p})$ approach $I_{p}(j_{p})$ as $j_{p}\to 0^{\mp}$.
$\lim_{j_{p}\to 0^{\mp}}I_{c|p}^{\pm}(j_{c}=0,j_{p})=I_{p}(j_{p}).$ (101)
Proof: For $j_{c}=0$, the saddle point equation (90) reduces to $\partial_{k}\hat{\omega}_{\mu_{\pm}}(k)|_{k={\rm i}\kappa_{\pm}}=0$, showing that $\kappa_{\pm}$ is independent of $j_{p}$. The result now follows directly from taking the limit $j_{p}\to 0^{\mp}$ in Eq. (100).
Claim 2: Derivatives of $I_{c,p}^{\pm}(j_{c}=0,j_{p})$ with respect to $j_{p}$ are finite and bounded by
<±$\mp\infty<\pm\partial_{j_{p}}I_{c,p}^{\pm}(j_{c}=0,j_{p}\in\mathcal{J}_{p}^{\mp})\leq\pm\partial_{j_{p}}I_{p}(j_{p}).$ (102)
@@ D.2 Generic properties of the split rate functions
$\partial_{j_{p}}I_{c,p}^{\pm}(j_{c},j_{p})=\partial_{j_{p}}I_{p}(j_{p})\pm\log\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm}),$ (103)
where reality of $I_{c,p}^{\pm}$ follows from (78). It is straightforward to show that $\hat{\omega}_{\mu_{\pm}}$ is strictly convex on the imaginary axis, $\partial_{x}^{2}\hat{\omega}_{\mu_{\pm}}({\rm i}x)=\int_{-\infty}^{\infty}{\rm d}\omega_{\mu_{\pm}}\,c^{2}e^{cx}>0$ while for $j_{c}=0$, the saddle point equation specifies that the derivative of $\hat{\omega}_{\mu_{\pm}}$ vanishes. Since the saddle points ${\rm i}\kappa_{\pm}$ are on the imaginary axis they are precisely the minima of $\hat{\omega}_{\mu_{\pm}}$ on the imaginary axis. Taking the value at the origin $\hat{\omega}_{\mu_{\pm}}(0)=1$, we have
@@
where finiteness follows from positivity of $\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})$. Plugging (104) into (103) gives the result.
Claim 3 : Derivatives of $I_{c,p}^{\pm}(j_{c}\neq 0,j_{p})$ with respect to $j_{p}$ diverge on inner boundaries of $\mathcal{J}_{p}^{\pm}$.
±$\pm\partial_{j_{p}}I_{c,p}^{\pm}(j_{c}\neq 0|j_{p}\to j_{p}^{\mp})\to\infty.$ (105)
@@
It follows that $\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})=\int_{-\infty}^{\infty}{\rm d}\omega_{\mu_{\pm}}(c)\,e^{\kappa_{\pm}c}\to\infty$ as $|\kappa_{\pm}|\to\infty$, which together with (103) gives the result.
Claim 4: The Hessian determinant of $I_{c,p}^{\pm}(j_{c},j_{p})$ is positive.
$\det{\rm H}_{I_{c,p}^{\pm}}(j_{c},j_{p})>0.$ (108)
@@
$\frac{{\rm d}}{{\rm d}u}\kappa_{\pm}(u)$ $=-\frac{\omega_{\mu_{\pm}}^{2}({\rm i}\kappa_{\pm})}{\partial_{k}^{2}\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})\omega_{\mu_{\pm}}({\rm i}\kappa_{\pm})+\left[\partial_{k}\hat{\omega}_{\mu_{\pm}}({\rm i}\kappa_{\pm})\right]^{2}}$
$=\frac{1}{h^{\prime}(\kappa_{\pm})}>0,$ (112)
@@
with positivity a consequence of (112) and strict convexity of $I_{p}$, giving the desired result.